### v1

# Optimizing Large Language Models for ESG Activity Detection in Financial Texts

Abstract
The integration of Environmental, Social, and Governance (ESG) factors into corporate decision-making is a fundamental aspect of sustainable finance. However, ensuring that business practices align with evolving regulatory frameworks remains a persistent challenge. AI-driven solutions for automatically assessing the alignment of sustainability reports and non-financial disclosures with specific ESG activities could greatly support this process. Yet, this task remains complex due to the limitations of general-purpose Large Language Models (LLMs) in domain-specific contexts and the scarcity of structured, high-quality datasets. In this paper, we investigate the ability of current-generation LLMs to identify text related to environmental activities. Furthermore, we demonstrate that their performance can be significantly enhanced through fine-tuning on a combination of original and synthetically generated data. To this end, we introduce ESG-Activities, a benchmark dataset containing 1,325 labeled text segments classified according to the EU ESG taxonomy. Our experimental results show that fine-tuning on ESG-Activities significantly enhances classification accuracy, with open models such as Llama 7B and Gemma 7B outperforming large proprietary solutions in specific configurations. These findings have important implications for financial analysts, policymakers, and AI researchers seeking to enhance ESG transparency and compliance through advanced natural language processing techniques.
Index Terms:
Deep learning, Environmental management, Financial technology, Generative AI, Large Language Models, Machine learning, Natural language processing, Sustainability, Text classification.
I Introduction
In recent years, driven by the widespread adoption of the Sustainable Development Goals (SDGs), the European Union has introduced principles and regulations aimed at helping organizations integrate environmental, social, and governance (ESG) factors into their operations and strategic decision-making. These initiatives encourage businesses and investors to assess and improve their environmental impact, fostering a more sustainable approach to economic activity [1]. As part of this effort, the ESG taxonomy [^1] https://finance.ec.europa.eu/sustainable-finance/tools-and-standards/eu-taxonomy-sustainable-activities_en has been established, offering a standardized framework to support sustainable practices across industries [2]. This resource enables companies to evaluate their activities in alignment with its criteria and report their performance in non-financial disclosures and sustainability reports.
However, aligning business practices with regulatory frameworks continues to be a challenge. Financial investors and other stakeholders often struggle to determine which aspects of corporate social responsibility correspond to specific activities regulated in the ESG taxonomy, making it difficult to accurately interpret and effectively utilize these reports. Indeed, a company can be associated with a vast amount of text that, in theory, can be mapped to various environmental activities described in the taxonomy. This text may include non-financial disclosures, marketing materials, website descriptions, product and service descriptions, corporate sustainability reports, and more. Thoroughly analyzing all this material is a complex and costly endeavour.
The rise of social fintech technologies, which increasingly support intelligent solutions for social impact investing, presents a promising opportunity to leverage advanced computational techniques for analyzing financial texts and extracting sustainability-related insights. For instance, several AI models have been developed to classify financial documents based on relevant SDGs [3, 4]. The rapid advancement of large language models (LLMs) in recent years has further improved automated solutions for the analysis of financial and sustainability-related texts [5, 6, 7, 8, 9]. However, categorizing text based on specific environmental activities, such as those defined in the ESG taxonomy, remains a significant challenge. Examples of these activities include promoting low-carbon rail transport, ensuring green port operations with zero direct CO2 emissions, or expanding infrastructure for cycling and personal electric mobility. This task is considerably more complex than linking text to broad SDG categories, as it requires a fine-grained understanding of the actions and activities described in a document. General-purpose LLMs struggle with this level of specificity. Overcoming these limitations, therefore, requires fine-tuning such models on high-quality, domain-specific datasets – an effort often hindered by the scarcity of training data in this domain.
In this paper, we explore the ability of current-generation LLMs to identify text related to environmental actions. Furthermore, we demonstrate how their performance can be significantly enhanced by fine-tuning them on a combination of original and synthetically generated data. Specifically, our study focuses on classifying textual segments extracted from Non-Financial Disclosures (NFDs) according to the activities defined in the ESG taxonomy. These activities include measures such as reducing emissions through energy-efficient production processes, transitioning to renewable energy, and implementing water conservation strategies, such as wastewater treatment and reuse in manufacturing plants.
To fine-tune and evaluate a range of models on their task, we introduce ESG-Activities, a novel benchmark consisting of 1,325 texts classified according to ESG activities. Since we intended assess the capability of synthetic data in supporting AI systems in this domain, the training set of ESG-Activities includes a mixture of manually curated data from human experts and synthetic data generated by LLMs. In contrast, the test set consists exclusively of items curated by human experts to ensure a high-quality evaluation.
The evaluation demonstrates that fine-tuning the system with a combination of manually crafted and synthetically generated data leads to significant improvements, not only compared to zero-shot learning but also over models fine-tuned exclusively on the manually annotated dataset. Furthermore, we show that relatively small open-source models, such as Llama 7B, can achieve excellent performance when fine-tuned on ESG-Activities, even surpassing large proprietary models in certain configurations.
In summary, the contributions of this paper are as follows:
  • We analyze the performance of state-of-the-art LLMs on the challenging task of associating textual descriptions with specific environmental activities.
  • We introduce ESG-Activities [^2] ESG-Activities - https://github.com/Mattia-Brt/Fine_tuning_LLM/tree/main/data , a novel benchmark that combines both original and synthetic data.
  • We demonstrate that augmenting a high-quality, manually curated dataset for fine-tuning with synthetic data enhances the performance of the resulting models on this task.
  • We provide the complete codebase for our analysis to ensure reproducibility [^3] Codebase - https://github.com/Mattia-Brt/Fine_tuning_LLM.
The remainder of this paper is structured as follows. Section II discusses the use case and provides a formal definition of the task. Section III introduces the new benchmark. Section IV describes the methodology adopted in our analysis. Section V presents the evaluation and discusses key insights. Section VI reviews related work. Finally, Section VII concludes the paper and outlines future research directions.
II Task definition and use case
II-A The role of ESG taxonomy for social fintech companies
In the European landscape, the analysis of social and environmental impact investments and the assessment of the sustainability of companies, including crowdfunding platforms or fintech, are becoming increasingly crucial. In 2021, the European Commission defined a set of ambitious policies related to the so-called “European Green Deal” with the overarching goal of making Europe climate-neutral by 2050. Among other initiatives, the European Green Deal has established a taxonomy that defines the activities companies must undertake to enhance their contributions to ESG aspects. This taxonomy is becoming a crucial benchmark for European investors, who utilize it to assess the sustainability of their investment decisions. Investors require a thorough understanding of the extent to which companies comply with the European ESG taxonomy. This is essential for making informed investment decisions that align with their values and contribute to sustainable development. However, the complexity of the ESG taxonomy and the vast amount of financial documentation can make it challenging for investors to independently assess a company’s compliance. As a result, there is an increasing need for tools that can streamline this process by automatically identifying and annotating relevant sections within financial documents. These tools represent sustainable applications of Social Fintech, showcasing how AI can enhance corporate social responsibility. By enabling investors to identify companies genuinely committed to sustainability, these technologies help direct capital toward initiatives aligned with SDGs, such as those outlined in the European Green Deal.
Furthermore, semi-automatic annotation tools can be seamlessly integrated into Social Fintech platforms, including crowdfunding platforms, reinforcing their dedication to sustainability and social impact. Automating ESG analysis with AI enhances transparency and promotes responsible investment opportunities, ultimately fostering a more sustainable and equitable financial system.
II-B Task definition
In this paper, we describe the process of developing an LLM tailored to verify if a text portion from NFD relates to a particular item in the ESG taxonomy. The task is framed as a binary classification problem, where the goal is to train the LLM to answer the question: Does c pertain to i?. Specifically, for a text segment $c$ and an item $i$ – in this case, the description of an ESG activity – the formulated fine-tuned LLM function $f(c,i)$ is expressed as:
$f(c,i)=\begin{cases}1&\text{ if c pertains to i},\\ 0&\text{otherwise}.\end{cases}$
The aim of the fine-tuned LLM is to learn a mapping function $f$ that accurately assigns each pair $(c,i)$ to the correct class.
This fine-tuning process is performed using a dataset of triples $(c,i,\text{class})$, where $c$ is a text chunk, $i$ is an activity description in the ESG taxonomy, and $\text{class}\in\{0,1\}$ indicates whether $c$ is related to $i$. The goal is to minimize the binary cross-entropy loss:
$\mathcal{L}=-\frac{1}{N}\sum_{j=1}^{N}\left[y_{j}\log\hat{y}_{j}+(1-y_{j})\log(1-\hat{y}_{j})\right],$
where $N$ is the number of samples in the dataset, $y_{j}$ is the true label (class) for the $j$-th sample, and $\hat{y}_{j}=f(c_{j},i_{j})$ is the predicted probability that $c_{j}$ pertains to $i_{j}$.
The fine-tuning process was performed using the Low-Rank Adaptation (LoRA) technique [10], which updates only a low-rank subspace of the model’s pre-trained weight matrices, reducing the number of trainable parameters while maintaining efficiency. Let $W$ be a weight matrix from the pre-trained model. During fine-tuning, the adaptation introduces a low-rank update $\Delta W$:
$\Delta W=AB,$
where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, $r\ll\min(m,n)$, ensuring efficiency in storage and computation.
The updated weight matrix becomes:
$W^{\prime}=W+\Delta W=W+AB.$
By freezing the original weights $W$ and only optimizing $A$ and $B$, LoRA enables efficient fine-tuning with minimal computational overhead.
III The ESG-Activities Benchmark
To the best of our knowledge, no existing dataset has been explicitly designed to enhance an LLM’s ability to classify text based on precise and granular environmental activities, such as those defined in the ESG taxonomy. To address this gap, we introduce ESG-Activities, a novel benchmark dataset. ESG-Activities links textual segments from NFDs to specific activities defined in the ESG taxonomy. The dataset was constructed following the process outlined below.
First, we selected the NFDs of four major companies within the transportation industry. The selected companies were Ferrovie dello Stato, the Italian State Railways; Autostrade per l’Italia, Italy’s largest company for motorway management and maintenance; Maersk, a well-known shipping and logistics company; and Mundys, a multinational company specializing in motorway and airport infrastructure. Next, we selected 12 activities related to the transport industry from the ESG taxonomy. These activities represent various actions that companies can take to promote sustainability in the transportation sector, such as developing zero-emission vessel infrastructure with electric charging and hydrogen refuelling, implementing smart mobility systems to enhance traffic efficiency, promoting low-carbon rail transport with CO2-free or bimodal trains, expanding infrastructure for walking, cycling, and personal electric mobility, and ensuring green port operations with zero direct CO2 emissions.
Since the original descriptions were too long, we generated shorter versions using GPT-4 as a rephraser. We then indexed the NFDs in a vector database and employed a Retrieval-Augmented Generation (RAG) [11] pipeline to retrieve the most relevant text chunks for each activity description. This process yielded 265 candidate mappings between text segments and ESG actions.
The candidate mappings were evaluated by three domain experts – professors and postdoctoral researchers with expertise in the transport industry. Each annotator independently assessed whether to confirm or reject each candidate mapping. A mapping was deemed valid if it received at least two positive votes out of three, following a majority rule. The resulting dataset consisted of 265 entries, each containing a textual description, an activity description, and a binary flag indicating a match. Among these, 78 text-activity pairs were classified as true matches. The dataset was then split into a training set of 212 instances and a test set of 53 instances.
Since we aimed to verify the effectiveness of syntactic data in this domain, we then expanded the training set with artificial data. Specifically, we used ChatGPT-4o to generate five alternative formulations for each of the original 212 sentences in the training set, ensuring they conveyed the same meaning while using different wording. This process resulted in 1,060 additional sentences, which were labelled as ’synthetic data’ and incorporated into the training set. This setup allows users of the benchmark to fine-tune their models using only the original data (212 items) or a combination of original and syntactic data (1272 items).
IV Methodology
In this section, we first introduce ESGQuest, the prototype we developed for analysing documents and mapping them to ESG activities using an LLM fine-tuned for this task. We then provide an overview of the LLMs evaluated in this study and describe the experimental setup used to assess their performance on the ESG-Activities benchmark.
IV-A System Architecture
The study presented in this paper is part of a broader project aimed at developing ESGQuest, a research prototype designed to support the annotation of NFD with ESG-related activities. This system is built on a RAG pipeline. Given a document about a company and a set of relevant NACE codes, ESGQuest aims to identify text segments that align with ESG activities associated with the specified NACE codes. The results are presented as an annotated PDF, which can be further refined by human users. [^4] A demo of this system is available at https://esgquest.datai.disco.unimib.it. The ESGQuest pipeline operates as follows.
First, ESGQuest selects ESG activities based on the given NACE codes using a predefined mapping schema. NACE codes (Nomenclature statistique des Activités économiques dans la Communauté Européenne) are a European industry standard classification system that groups organizations based on their economic activities.555 These codes are used in the EU Taxonomy to assess the environmental sustainability of economic activities. For more details, visit https://ec.europa.eu/eurostat/web/nace This step is optional but generally beneficial, as many ESG activities described in the taxonomy are relevant only to specific industrial sectors (e.g., transport, energy). Including irrelevant activities in the analysis would not only waste computational resources but also introduce unnecessary complexity and potential confusion.
Second, the input documents are divided into smaller chunks and stored in Pinecone [^6] Pinecone - https://www.pinecone.io/, a vector database optimized for similarity search. Pinecone facilitates the efficient retrieval of the most semantically relevant segments for a given query by leveraging high-dimensional vector embeddings. It is specifically designed for approximate nearest neighbour search, ensuring fast and scalable retrieval even for large datasets.
Third, for each relevant ESG activity, the system queries the vector database with the activity description to retrieve the most pertinent chunks from the input documents.
Finally, an LLM optimized for this purpose analyzed each chunk to assess whether it aligned with the description of the action. If a chunk is assigned to an action, the corresponding text segment in the original PDF is annotated accordingly.
IV-B Large Language Models
In this paper, we examine a range of LLMs, including several open-source models, which offer a cost-effective solution and can be deployed locally to ensure the privacy of sensitive documents, as well as a large proprietary model (GPT-4o Mini). Below, we provide a description of each LLM evaluated in this paper, including their sizes and key characteristics.
IV-B1 Llama.
Llama, introduced by Hugo Touvron et al. in 2023 [5], is a family of decoder-only language models designed to deliver high performance and computational efficiency. The Llama models range from 1 to 90 billion parameters, making them suitable for a wide variety of natural language processing applications. In this study, we evaluate three versions: Llama 3B, Llama 2 7B, and Llama 3 8B, all available on official platforms such as Hugging Face.
Llama 3B [^7] https://huggingface.co/openlm-research/open_Llama_3b is an open-source model with 3 billion parameters. The model incorporates architectural enhancements such as pre-normalization and the SwiGLU activation function [12], which contribute to training stability and efficiency.
Llama 2 7B [^8] https://huggingface.co/meta-Llama/Llama-2-7b-hf, represents an evolution of the first generation of Llama [13]. It was trained using additional data and optimized techniques to improve generalization. This model excels in zero-shot and few-shot learning tasks and directly competes with larger models like GPT-3 and Chinchilla [14].
Llama 3 8B [^9] https://huggingface.co/meta-llama/Meta-Llama-3-8B, is the latest iteration in the series. This model, featuring 8 billion parameters, was trained on a dataset exceeding 15 trillion tokens [15]. It incorporates significant advancements such as odds ratio preference optimization [16], which optimizes performance and alignment.
The Llama series was trained on diverse datasets, comprising publicly available data such as CommonCrawl, C4 [17], GitHub, Wikipedia, Books3 [18], ArXiv, and Stack Exchange.
These models have demonstrated excellent performance in benchmarks such as NaturalQuestions [19] and TriviaQA [20], often surpassing larger models like Gopher (280B) [21] and Chinchilla (70B) [14]. For instance, the Llama 33B model outperformed both Gopher and Chinchilla in question-answering tasks on the NaturalQuestions benchmark, demonstrating that the Llama architecture effectively balances size and performance.
IV-B2 Gemma.
Introduced by Google in 2024, Gemma [6] is a family of open-source language models derived from the Gemini series [7], designed to achieve an optimal balance between computational efficiency and performance across various natural language processing tasks. The models are available in three configurations – 2B, 7B, and 27B parameters – each tailored for specific applications. Gemma employs a decoder-only architecture with advanced features such as multi-query attention to reduce memory overhead [22], Rotary Position Embeddings (RoPE) [17] to improve positional representation, GeGLU activations for better convergence during training, and RMSNorm to enhance stability on large-scale datasets. The models were trained on diverse publicly available data, including CommonCrawl, C4 [17], GitHub repositories, ArXiv papers, Wikipedia, and Books3 [18], ensuring broad linguistic coverage and domain-specific expertise. The 2B model [^10] https://huggingface.co/google/gemma-2b is optimized for on-device deployments and CPU-based inference, making it efficient for general-purpose applications. The 7B variant [^11] https://huggingface.co/google/gemma-7b is fine-tuned for complex tasks like advanced coding, scientific reasoning, and mathematical problem-solving, outperforming comparable models such as Llama 7B [5] and often rivalling Mistral 7B [8]. The Recurrent Gemma 2B model [^12] https://huggingface.co/google/recurrentgemma-2b introduces a recurrent mechanism that enhances performance on sequential reasoning tasks and long-context processing, making it particularly effective for multi-turn dialogues and structured data analysis [23]. Across multiple benchmarks, including coding, reasoning, and question answering, Gemma models have demonstrated competitive performance.
IV-B3 Mistral.
Mistral, introduced by Albert Q. Jiang et al. in 2023 [8], is a highly efficient decoder-only language model designed for a wide array of natural language processing tasks, excelling in areas such as reasoning, mathematics, code generation, and commonsense reasoning. With 7.3 billion parameters, Mistral was specifically engineered to achieve a balance between computational efficiency and performance. The model incorporates several advanced techniques that make it highly effective for tasks that demand long-context processing and memory optimization. One such technique is the sliding window attention (SWA) [24], which enables the model to process long sequences of input in overlapping chunks, allowing it to handle inputs of arbitrary length with minimal computational overhead. In addition, Mistral utilizes grouped-query attention (GQA) [25], a method that optimizes inference speed by reducing memory usage during the attention computation phase.
Mistral’s ability to maintain performance despite having a smaller parameter count than other models makes it a powerful tool for practical applications, especially where computational resources may be limited. It consistently outperforms larger models like Llama 2 13B [5], Llama 1 34B [26], and other established benchmarks across a variety of metrics. This is especially evident in reasoning-intensive tasks, where Mistral exhibits superior generalization capabilities.
The 7B variant of Mistral [^13] https://huggingface.co/mistralai/Mistral-7B-v0.1 has been fine-tuned to support diverse use cases, including scientific research, where accurate interpretation and generation of complex texts are critical. The model is also capable of achieving state-of-the-art results on tasks requiring a deep understanding of specialized knowledge, such as those found in academic papers or technical documents.
IV-B4 Gpt-4o Mini.
GPT-4o Mini is a decoder-only language model developed by OpenAI as a faster and more cost-effective alternative to GPT-4o. GPT-4o is a comprehensive multimodal model capable of processing and generating text, images, and audio, designed for complex tasks requiring high computational power and precision. It features a 128K-token context window and supports up to 16K output tokens per request, with a knowledge cutoff in October 2023 [9]. Despite its smaller size, GPT-4o Mini outperforms previous models, such as GPT-3.5 Turbo, in academic benchmarks, demonstrating superior performance in textual intelligence and multimodal reasoning [27]. It retains the same context window and output token limits as GPT-4o, ensuring robust performance across various applications. Additionally, GPT-4o Mini incorporates an instruction hierarchy method to enhance resistance to jailbreaks and prompt injections, improving security for deployment in diverse settings [28].
IV-C Experimental Setup
In order to study the advantages of fine-tuning LLMs for ESG activity classification, we evaluated the models in two different modes: zero-shot and fine-tuning. The fine-tuning was performed on the two alternative training sets within ESG-activities: the manually curated dataset of 212 instances and a larger dataset of 1,272 instances, which combines original and synthetic data. In the following sections, we provide a detailed discussion of both approaches.
IV-C1 Zero-shot
Zero-shot learning is a method in which a LLM is prompted to perform a task based solely on provided instructions, without the need for task-specific examples. This approach leverages the extensive pretraining of LLMs, allowing them to generalize to new, previously unseen tasks. It is particularly valuable because it eliminates the need for additional task-specific training data, making it a cost-effective and scalable solution across various applications.
In this study, we designed and iteratively refined a prompt template for zero-shot classification. The template takes two inputs: 1) the text to be classified and 2) a description of an ESG action. Based on these inputs, the prompt instructs the LLM to return a value of 1 if the text aligns with the ESG action and 0 otherwise.
IV-C2 Fine-tuning
Fine-tuning is a supervised learning technique that involves adjusting the weights of a pre-trained LLM using a labelled dataset. This process refines the model’s broad language understanding and knowledge acquired during its initial pre-training phase, enabling it to achieve superior performance on specific tasks or within particular domains.
In this study, we fine-tuned eight LLMs for the proposed task using the training and validation datasets from the ESG-Action benchmark. Specifically, each model was fine-tuned twice: first using the manually curated training and validation datasets, and then using augmented datasets that included synthetic data. This approach enables a comparative analysis of the two model versions, allowing us to evaluate the impact of synthetic data on models of different sizes.
Due to the limited size of our training dataset, we employed a 10-fold cross-validation approach to mitigate the risk of overfitting. This technique systematically partitions the dataset into ten subsets, ensuring that the model is trained and evaluated on different data portions in each iteration. By rotating the validation set across all folds, cross-validation reduces bias and variance, leading to a more reliable assessment of the model’s generalization capability and performance metrics. Models were then saved after each epoch, and validation was conducted at regular intervals, referred to as steps. Specifically, the models were evaluated and saved every 10% of the total optimization steps. This strategy enables the selection of the best-performing model based on validation metrics and mitigates issues like overfitting or underfitting.
The optimization process was handled using the AdamW optimizer [29], a variant of the Adam optimizer that incorporates decoupled weight decay regularization. To optimize the fine-tuning process, we employed LoRA techniques from the PEFT library, as previously mentioned. This approach reduces the number of trainable parameters while preserving model performance. For instance, when fine-tuning Llama 3B, which has over 3.3 billion parameters, LoRA required modifying only 0.08% of the total parameter space. Similarly, for Llama 7B, only 0.12% of its parameters were updated.
The only exception to this procedure was GPT-4o Mini, which we fine-tuned using Azure OpenAI [^14] Azure OpenAI - https://azure.microsoft.com/en-us/products/ai-services/openai-service, a cloud-based service provided by Microsoft that enables access to OpenAI’s language models. We formatted the training dataset in the required JSON structure, uploaded it to Azure, and executed the fine-tuning process with the default hyperparameters.
V Evaluation
In this section, we present the results of our experiments on the ESG-Activities benchmark. Specifically, we evaluated the eight LLMs described in Section IV using three alternative configurations: 1) zero-shot learning, 2) fine-tuning on the original data from ESG-Activities, and 3) fine-tuning on a combination of the original and synthetic data. All models were tested on the test set of ESG-Activities.
As the task is framed as a binary classification problem, we used precision, recall, and F1 score as evaluation metrics.
V-A Zero-Shot Setting
Table I reports the performance of the models in a zero-shot setting. GPT4o_mini yields the best result (72.7%), followed by GEMMA_7B (70.5%), and Llama3_8B (70.1%). GEMMA_2B and Llama_3B performed fairly well, exceeding 68% F1. Conversely, RecurrentGEMMA_2B and MISTRAL showed relatively poor performance in this task. This suggests that the recurrent architecture of RecurrentGEMMA_2B may not offer significant benefits in this specific task. Overall, there is considerable room for improvement.
Model Precision Recall F1-Score
Llama_3B 0.6565 0.7358 0.6816
Llama2_7B 0.7145 0.5472 0.5828
Llama3_8B 0.6845 0.7358 0.7008
GEMMA_2B 0.6696 0.6792 0.6742
GEMMA_7B 0.8361 0.6792 0.7050
recurrentGEMMA_2B 0.7455 0.4717 0.4945
MISTRAL 0.5574 0.3207 0.3396
GPT4o_mini 0.7202 0.7358 0.7270
TABLE I: Performance of the models in a zero-shot setting.
V-B Fine-tuning
Table II presents the results of the models fine-tuned on the manually curated dataset. The results demonstrate a significant improvement across all models, confirming the efficacy of the ESG-Action benchmark in optimizing performance for this task.
The top-performing model is again GPT4o_mini, which attains an F1 score of 80.5%, followed by Llama_3B (78.9%) and GEMMA_2B (78.1%). It is interesting that the smaller model achieved the best performance after fine-tuning. This is likely due to the fact that the manually curated training set is relatively small. As a result, models tend to overfit quickly, and this effect is even more pronounced in larger models, which overfit earlier and thus experience a drop in performance.
Model Precision Recall F1-Score
Llama_3B 0.7866 0.7925 0.7892
Llama2_7B 0.8361 0.6792 0.7050
Llama3_8B 0.7420 0.7170 0.7273
GEMMA_2B 0.7763 0.7925 0.7813
GEMMA_7B 0.8146 0.6981 0.7226
recurrentGEMMA_2B 0.7763 0.4340 0.4396
MISTRAL 0.6701 0.5472 0.5838
GPT4o_mini 0.8015 0.8113 0.8050
TABLE II: Performance of the fine-tuned models (original data).
Table III presents the performance of models fine-tuned on the second training set of ESG-activities, which includes a combination of original and synthetic data. This experiment aimed to determine whether incorporating synthetic data during fine-tuning provides a performance advantage. The results clearly indicate that it does, with most models now achieving F1 scores exceeding 80%. The best overall result is achieved by Llama2_7B, with an F1 score of 84.9%, followed by Llama_3B at 84.4% and GEMMA_7B at 82.1%. GPT4o_mini, which was fine-tuned using the OpenAI service, did not perform as well in this setting.
Model Precision Recall F1-Score
Llama_3B 0.8421 0.8491 0.8440
Llama2_7B 0.8491 0.8491 0.8491
Llama3_8B 0.8290 0.7925 0.8036
GEMMA_2B 0.8176 0.8302 0.8127
GEMMA_7B 0.8190 0.8302 0.8211
recurrentGEMMA_2B 0.8470 0.5283 0.5452
MISTRAL 0.8032 0.7736 0.7840
GPT4o_mini 0.7684 0.7925 0.7711
TABLE III: Performance of the fine-tuned models (original and synthetic data).
In conclusion, our experiments demonstrate that LLMs can achieve high performance in classifying text based on ESG activities when fine-tuned on high-quality datasets. These findings also provide further evidence supporting the effectiveness of synthetic data in specific domains. Notably, the most comprehensive version of the ESG-Activities benchmark training set, which combines both manually annotated and synthetic data, further improves classification performance. For example, Llama2_7B, the best-performing model, achieved a remarkable 14.8% increase in F1 score, rising from 70.1% to 84.9% with the inclusion of synthetic data.
V-C Training and Inference Time
Training times varied significantly across models, reflecting differences in architecture, size, and the complexity of the fine-tuning processes. Analyzing these variations offers valuable insights into the computational demands of each model.
Table IV provides the training and inference times for the various models. It was not possible to report the fine-tuning time for GPT4o_mini, as this process was managed by the OpenAI service.
Model Inference Time Training Time Training Time
(s) Original (s) Original+Synthetic (s)
Gemma_2B 3.24 115 828
Gemma_7B 2.18 120 976
RecGemma_2B 2.51 130 1011
Llama3_8B 3.74 226 1096
Llama2_7B 2.14 132 1046
Llama_3B 3.61 181 1534
Mistral_7B 10.89 703 5888
GPT4o_mini 4.46 NA NA
TABLE IV: Training and inference time for the eight models.
Notably, models from the Gemma family exhibit faster training times compared to the Llama models. Among the models analyzed, MISTRAL_7B stands out with a significantly longer training time of about 96 minutes, nearly seven times that of GEMMA_2B. Regarding inference times, Llama2_7B is the fastest model, followed by Gemma_7B and RecurrentGemma_2B.
V-D Cost Analysis
The computational demands of training and inference processes can have a significant impact on the overall cost of a project.
Model original orig+syntetic
GPT4o_mini 0.14 0.86
Gemma_2B 0.29 2.07
Gemma_7B 0.30 2.44
RecGemma_2B 0.33 2.52
Llama2_7B 0.33 2.62
Llama3_8B 0.56 2.74
Llama_3B 0.45 3.84
Mistral_7B 1.76 14.72
TABLE V: Fine-tuning costs
Table V reports the costs associated with fine-tuning the eight models on both the original dataset and the full version, which includes synthetic data. All models, except GPT4o_mini, were trained on an NVIDIA A100 GPU. The hourly rental rate for an A100 machine was approximately \$9 at the time of this study. Differently, the fine-tuning of GPT4o_mini was conducted using OpenAI’s standard API offering, which, at the time of writing, allowed fine-tuning at a cost of \$3.00 per 1 million tokens. For these reasons, although GPT4o_mini is the largest model in terms of the number of parameters, it is the cheapest to train using OpenAI’s services. However, this training cost must be combined with the fees required to deploy and maintain the model on OpenAI’s infrastructure, which is currently \$1.62 an hour. Among the open models, the most cost-effective to fine-tune with our configuration are Gemma_2B and Gemma_7B.
VI Related work
This section examines key advancements in the development and application of LLMs in the ESG domain. LLMs have demonstrated remarkable capabilities in processing textual data at scale, enabling the accurate extraction and synthesis of information from diverse sources, including research papers [30], patents [31], medical records [32], social media posts [33], news articles [34], technical documentation [35], legal texts [36], and financial documents [37, 38, 39].
Recent LLM families, including GPT [9], Llama [5], Mistral [8], and Gemma [6], have demonstrated outstanding capabilities in analyzing and classifying ESG and financial data. Their proficiency in processing vast amounts of unstructured information positions them as powerful tools for enhancing ESG reporting, regulatory compliance, and data-driven decision-making. However, the standard versions of these models, when used with simple prompting, do not always achieve optimal performance. Consequently, many researchers focus on developing domain-specific adaptations by fine-tuning open models on ESG and financial data. For instance, Yang et al. [40] introduced FinGPT, an open-source LLM designed for the financial sector. Initially trained on a general-purpose corpus, FinGPT was then fine-tuned using domain-specific datasets, including financial news, earnings reports, and real-time market data. This process enables the model to capture the nuanced complexities of financial language, improving its performance in tasks such as market forecasting, sentiment analysis, and news classification. As a result, FinGPT can be applied to various use cases, including robo-advising, algorithmic trading, and low-code development.
FinBERT [41] is a domain-specific adaptation of BERT [42] designed for financial text analysis. It is pre-trained on a vast corpus of financial documents, including regulatory filings, corporate announcements, and financial articles, before undergoing fine-tuning for specific tasks. This specialized training enables FinBERT to develop a deep understanding of financial terminology and contextual nuances. Notably, it has demonstrated exceptional performance in sentiment classification of financial analyst reports. Similarly, ESG-BERT [43] is a specialized variant of BERT that has been pre-trained on a large corpus of sustainability-related texts. This model has demonstrated strong performance in various natural language processing tasks within the ESG domain, particularly in text classification and sentiment analysis [44]. We initially explored adapting these models to the task analyzed in this paper; however, this approach proved unfeasible. These models are trained for specific classification tasks with predefined labels, making it difficult to adapt them for matching a text with an arbitrary activity description.
Several approaches have been proposed for ESG sentiment classification, which involves categorizing text as either positive/opportunity or negative/risk from an ESG perspective. Notably, the ML-ESG-2 shared task [45], co-located with the FinNLP workshop at IJCNLP-AACL 2023, has played a key role in advancing methodologies in this field. To tackle this challenge, researchers have leveraged both encoder-only and decoder-only transformer models, including fine-tuned versions of FinBERT [46] and LLMs such as Llama 2 and Dolly  [47]. Similar to previous approaches, we leverage and evaluate a variety of fine-tuned language models. However, the task addressed in this paper is more fine-grained than merely detecting a positive or negative sentiment.
To tackle the challenges of long contexts and data recency in LLMs, researchers have started incorporating the RAG paradigm [11], which improves response accuracy by dynamically retrieving relevant information. For instance, ESGReveal [47] is a recent method for extracting and analyzing ESG data from corporate reports that utilizes an LLM augmented with RAG. The system includes an ESG metadata module for precise querying, a preprocessing module for database construction, and an LLM agent for extracting information. Like our approach, ESGReveal leverages an RAG architecture. However, its primary focus is on extracting quantitative data from documents, such as Scope 1 and Scope 2 emissions. In contrast, our work is centred on identifying portions of documents that align with the activities described in the ESG taxonomy. Furthermore, ESGReveal utilizes zero-shot learning LLMs with basic prompting, whereas we fine-tune LLMs to optimize their performance for our specific task.
Efforts have also been made to develop a question-answering system in the ESG domain using a pipeline based on Knowledge Graph-based Retrieval Augmented Generation (KG-RAG) [48]. Unlike traditional RAG, which retrieves information from textual sources, KG-RAG leverages knowledge graphs [49], which are structured representations of information that capture relationships between entities in a graph format. Similarly, Angioni et al. [50] employ a combination of linguistic pattern analysis and transformer models to extract a knowledge graph from a corpus of news articles, aiming to capture key trends related to ESG aspects. In this work, we focus on the more specific task of classifying text based on ESG activities. Consequently, we rely exclusively on action descriptions from the ESG taxonomy rather than employing a more complex knowledge representation.
VII Concusion
In this paper, we investigate the ability of current-generation LLMs to identify text related to ESG activities. Specifically, we assess strategies for optimizing these models using high-quality data and explore the feasibility of leveraging a combination of original and synthetically generated data. To support this study, we introduce ESG-Activities, a novel benchmark comprising 1,325 labelled text segments classified according to the EU ESG taxonomy.
Our results demonstrate that fine-tuned models significantly outperform zero-shot approaches, confirming that domain adaptation plays a crucial role in enhancing ESG text classification. Moreover, the inclusion of synthetic data proved to be a valuable strategy, yielding substantial performance gains across multiple model architectures. This finding suggests that synthetic data augmentation can serve as an effective method to overcome data scarcity issues, particularly in specialized domains such as ESG analysis.
Another key takeaway from our experiments is that relatively small, open-source models can achieve competitive performance levels when fine-tuned with high-quality, domain-specific data. This insight has significant implications for organizations looking to deploy cost-effective AI solutions for ESG analysis or those that need to run models locally due to privacy and data security concerns, which are particularly common in the financial domain.
This study opens several promising avenues for future research. First, we plan to explore alternative data generation techniques, including human-in-the-loop approaches, to further refine synthetic data for ESG applications. Second, while our work focused on ESG-related sentence classification, further research is needed to evaluate the models’ performance on other NLP tasks, such as ESG sentiment analysis and risk assessment. Finally, future studies should examine how these models can be effectively deployed in real-world financial decision-making processes, ensuring that AI-driven ESG analysis aligns with regulatory requirements and industry standards.
Acknowledgment
This work is supported by the Italian Ministry of University and Research (MUR) within the PRIN2022—ISALDI: Interpretable Stock Analysis Leveraging Deep multImodal models (CUP: E53D23008150006).